(no subject)

arenton@uvahef.phys.virginia.edu
Fri, 30 Jul 1999 08:01:23 -0400

Shift Summary for owl shift, 30/july/99

On Shift: Mike Arenton, Jason LaDue

Hours of beam delivered at all intensities: 7
Hours of beam delivered at nominal intensity: 7
Hours of E832 data (special runs not included): 6.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: tape change, kfrend's etc.

========================================================================
Runs ended this shift:
Run 14061 BEAM832 9.2M events, 13385 vac 00's

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
very smooth E832 datataking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14061 02:02 run ends as tapes full

14062 02:20 new E832 run after DART fresh start

14062 65-72 03:50 AZLAT DYC full, Kfrend doesn't work,
power cycle downstairs repeater fixes it.

14062 100-101 04:33 ktevana errors, kfrend.

14062 200-203 06:44 CsI 1297 has CapId problem, and is not
in laser illumination. Pipeline reset
appears to fix the problem.

========================================================================
Follow up Notes on Previous Shift

Still getting red SDAQ button on Etot-temp

========================================================================
Items to Follow up on Next Shift

Etot-temp should be looked at by experts.